### v2

# Gagan Bhatia ξ El Moatez Billah Nagoudi ξ Abdellah El Mekki ω Fakhraddin Alwajih ξ Muhammad Abdul-Mageedξ,ω,λ

Abstract
We introduce Swan, a family of embedding models centred around the Arabic language, addressing both small-scale and large-scale use cases. Swan includes two variants: Swan-Small, based on ARBERTv2, and Swan-Large, built on ArMistral, a pretrained Arabic large language model. To evaluate these models, we propose ArabicMTEB, a comprehensive benchmark suite that assesses cross-lingual, multi-dialectal, multi-domain, and multi-cultural Arabic text embedding performance, covering eight diverse tasks and spanning 94 datasets. Swan-Large achieves state-of-the-art results, outperforming Multilingual-E5-large in most Arabic tasks, while the Swan-Small consistently surpasses Multilingual-E5-base. Our extensive evaluations demonstrate that Swan models are both dialectally and culturally aware, excelling across various Arabic domains while offering significant monetary efficiency. This work significantly advances the field of Arabic language modelling and provides valuable resources for future research and applications in Arabic natural language processing. Our models and benchmark will be made publicly accessible for research.
[Uncaptioned image] Swan and ArabicMTEB: Dialect-Aware, Arabic-Centric, Cross-Lingual, and Cross-Cultural Embedding Models and Benchmarks
Gagan Bhatia ξ El Moatez Billah Nagoudi ξ Abdellah El Mekki ω Fakhraddin Alwajih ξ Muhammad Abdul-Mageedξ,ω,λ ξ The University of British Columbia,      ω MBZUAI,      λ Invertible AI {gagan30@student.,muhammad.mageed@}ubc.ca
1 Introduction
Natural Language Processing (NLP) has seen rapid advancements in recent years, driven by groundbreaking developments in deep learning and the emergence of sophisticated distributed text representations such as word and sentence embeddings Devlin et al. (2018); Reimers and Gurevych (2019). These embeddings, which transform text into dense vectors, enable effective semantic understanding and are pivotal for enhancing performance across many downstream applications, including text classification, semantic search, and machine translation. Moreover, text embeddings have become fundamental to the success of large language models (LLMs) Touvron et al. (2023b, a); Jiang et al. (2023); Team et al. (2024), which are increasingly integrated into a variety of real-world systems and tools. One of the most promising applications of these embeddings is in the realm of Retrieval-Augmented Generation (RAG) Shao et al. (2023); rag (2023), where LLMs are augmented with information retrieval capabilities. In RAG-based systems, lightweight embedding models retrieve relevant information from large corpora, fed as context to models like ChatGPT OpenAI (2023) or GPT-4 OpenAI et al. (2024). This synergy between embeddings and LLMs has demonstrated significant improvements in both general-purpose tasks such as question answering Lin et al. (2023); rag (2023) and domain-specific applications Bhatia et al. (2024); Shi et al. (2023); Lin et al. (2023).
Figure 1: Details of ArabicMTEB
Despite these advances, the predominant focus of current embedding models has been on English and Chinese, which limits their applicability to other languages. This is particularly true for Arabic, a language with rich morphology, diverse dialects, and unique syntactic structures, making it challenging to develop effective language representations Nagoudi et al. (2022); Huang et al. (2024). Existing multilingual models often fail to capture these complexities, leading to a suboptimal performance on Arabic NLP tasks Abdul-Mageed et al. (2020a); Elmadany et al. (2022). Addressing this limitation requires the development of Arabic-specific embedding models that are sensitive to the linguistic and cultural nuances of the language.
In this work, we introduce Swan, a family of dialect-aware Arabic-centric cross-lingual cross-cultural embedding models designed to bridge this gap and push the boundaries of Arabic NLP. Our contributions are as follows:(1) We introduce Swan, a cutting-edge family of Arabic embedding models. This includes two variants: Swan-Small, based on ARBERTv2 Elmadany et al. (2022), and Swan-Large, built upon ArMistral, a further pretrained Arabic language model. (2) We present ArabicMTEB, a comprehensive evaluation benchmark for Arabic text. ArabicMTEB is designed to assess cross-lingual, multi-dialectal, multi-domain, and multi-cultural performance, spanning eight tasks and $94$ datasets. (3) Our larger model, Swan-Large, showcases state-of-the-art text embedding capabilities, surpassing Multilingual-E5-large Wang et al. (2024b) on most Arabic tasks. Additionally, our smaller, Swan-Small, consistently outperforms Multilingual-E5-base Wang et al. (2024b) on most Arabic tasks. (4) Through rigorous benchmarking, we demonstrate that Swan models are not only dialectally and culturally aware but also excel across diverse Arabic domains while maintaining a significantly lower monetary cost.
The rest of the paper is organized as follows: In Section 2, we review related work with a particular emphasis on Arabic text embedding models, their applications and challenges. We present our approach to model training Swan models in Section 3. Section 4 outlines how we built our benchmark dataset, ArabicMTEB. Section 5 is about our experiments and model analysis. Finally, we conclude in Section 6.
2 Related Work
Benchmark Language Tasks Datasets #Tasks CRTR Arabic Culture/Domains
MTEB Muennighoff et al. (2022) English RTR, STS, PairCLF, CLF, RRK, CLR, SUM 56 7 $\times$ $\times$
C-MTEB Xiao et al. (2023) Chinese RTR, STS, PairCLF, CLF, RRK, CLR 35 6 $\times$ $\times$
De-MTEB Sturua et al. (2024) German RTR, STS, PairCLF, CLF, RRK, CLR 17 6 $\times$ $\times$
F-MTEB Ciancone et al. (2024) French RTR, STS, PairCLF, CLF, RRK, CLR, BTM 17 7 $\times$ $\times$
Es-MTEBMohr et al. (2024) Spanish RTR, STS, PairCLF, CLF, RRK, CLR 17 6 $\times$ $\times$
Polish-MTEB Poświata et al. (2024) Polish RTR, STS, PairCLF, CLF, CLR 26 5 $\times$ $\times$
Ru-MTEB Poświata et al. (2024) Russian RTR, STS, PairCLF, CLF, RRK, CLR 23 6 $\times$ $\times$
Scand. MTEB Enevoldsen et al. (2024) Danish RTR, CLF, BTM, CLR 26 4 $\times$ $\times$
Norwegian $\times$ $\times$
Swedish $\times$ $\times$
ArabicMTEB (Ours) Arabic RTR, STS, PairCLF, CLF, RRK, CLR, BTM, CRTR 94 8
Table 1: Comparison of Various Text Embedding benchmarks proposed in the literature across the different covered task clusters. RTR: Retrieval, STS: Semantic Textual Similarity, PairCLF: Pair Classification, CLF: Classification, CLR: Clustering, RRK: Reranking, BTM: BitextMining, CRTR: Crosslingual Retrieval.
In recent years, there have been remarkable advancements in text embedding models, with a shift towards developing universal embeddings for diverse tasks and domains. Despite this, specialized models and benchmarks for languages like Arabic remain underexplored.
Multilingual Text Embeddings Models. With the need for language-agnostic embeddings growing, multilingual models like LASER Artetxe and Schwenk (2019) and LaBSE Feng et al. (2022) were developed using BiLSTM and Transformer encoders, respectively. Building on this, the Multilingual-E5 Wang et al. (2024c) series extends the E5 architecture to support diverse languages using multilingual text pairs and synthetic data. GRIT Muennighoff et al. (2024) further unifies generative and embedding tasks within a single model. Newer models like ColBERT-XM Louis et al. (2024) and Gecko Lee et al. (2024) refine multilingual embeddings through modular and distilled architectures.
Arabic-Specific Models. Despite progress in Arabic NLP, models are not optimized for Arabic text embedding and retrieval. Efforts like ARBERT Abdul-Mageed et al. (2021a), ORCA Elmadany et al. (2023), and AraMus Alghamdi et al. (2023) have focused on encoding and generation but are not tailored for sentence-level embeddings. While language-agnostic models such as LASER and Multilingual-E5 include Arabic in their training data, they may not fully capture its linguistic intricacies. To address this, Nacar and Koubaa (2024) introduced models and training datasets to improve semantic similarity performance for Arabic.
Text Embedding Benchmarks. Most text embedding evaluations rely on a narrow set of datasets, limiting their generalisation ability. To address this, the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023) introduced eight task categories with 58 datasets and 112 languages. However, it remains predominantly focused on English. Similar benchmarks have been developed for other languages, such as C-MTEB Xiao et al. (2023) for Chinese. For Arabic, evaluations have primarily centred on Semantic Text Similarity (STS) tasks Nacar and Koubaa (2024). However, excelling in STS does not guarantee optimal performance in tasks like clustering or reranking Muennighoff et al. (2023). Existing Arabic benchmarks like ORCA Elmadany et al. (2023) and ALUE Seelawi et al. (2021) focus on natural language understanding (NLU), while Dolphin Nagoudi et al. (2023a) targets natural language generation (NLG). This work is the first comprehensive benchmark for evaluating Arabic text embeddings across multiple tasks.
3 Swan
3.1 Training Data
Family Language Type Dataset Level Size
Monolingual Arabic Human ORCA-MSA Sentence 378K
ORCA-DIA 122K
MMARCO-ar 8.1M
Synthetic Synth-MSA Paragraph 100K
Synth-DIA 15K
Synth-DOM 20K
Crosslingual Arabic to 15 Langs Human MMARCO Sentence 3M
Arabic to 6 Langs XOR-TyDi 20.5K
Multilingual 11 Langs Human Mr-Tydi Sentence 49K
16 Langs Miracl 343K
Total 12.5M
Table 2: The diverse datasets employed for training our Arabic embedding models. In the synthetic dataset, we have three datasets: the MSA dataset, the Dialectal dataset (Egyptian and Moroccan), and domain-based focusing on Medical, Financial, Legal and News domains.
To train Swan, we develop the most extensive training corpus for Arabic embedding models, leveraging a unique assembly of datasets to ensure comprehensive linguistic coverage and diversity. Our training data covers paragraph-based and sentence-based datasets curated from multiple sources. Table 2 shows an overview of our training datasets.
You have been assigned a retrieval task: {task} Your mission is to write one text retrieval example for this task in JSON format. The JSON object must contain the following keys: new user_query: a string, a query specified by the retrieval task. new positive: a string, a relevant document for the user query. new hard_negative: a string, a document closely related to the query. Please adhere to the following guidelines: The user_query should be paragraph-based, understandable with some effort or ambiguity, and diverse in topic. The hard_negative contains some useful information, but it should be less useful or comprehensive than the positive. Refer to caption Refer to captionCohere
(a) Positive and hard negative generation
Figure 2: Methodology to generate our synthetic data.
MSA Datasets. We focus on two sources: (i) Human-generated data: Composed from ORCA Elmadany et al. (2023) and mMARCO Bonifacio et al. (2021). ORCA is a compilation of labelled datasets with tasks such as semantic text similarity (STS), sentence classification, text classification, natural language inference (NLI), and question answering. We use all the training sets from ORCA, encompassing $60$ different datasets. mMARCO-ar is the translated version of MARCO, which is a human-generated dataset Bajaj et al. (2018). Both of these datasets are cleaned up and de-duplicated using Polydedupe  Bhatia (2023), which is further described in Appendix D. (ii) Synthetically-generated data: To augment our MSA training data for retrieval tasks, we use Command R+ Cohere For AI (2024) to generate high-quality synthetic data. [^1] We performed various in-house evaluations comparing multiple models. Command R+ was chosen as it is open-source and efficient in generating Arabic varieties (standard and dialectal). The generation methodology was inspired by Wang et al. (2024a). We first generate tasks closely related to real-world usage using the model, and then we use the procedure shown in Figure 2 to generate our synthetic dataset. We generate $100k$ in general MSA data and $5k$ in instances for specific domains like finance, news, medicine, and legal for a total of $120k$ MSA instances.
Dialectal Arabic Datasets. Similar to the MSA datasets, we focus on two sources: (i) Human-generated data: We use publicly available dialectal Arabic data, which primarily covers Gulf, Egyptian, Moroccan, and Levantine varieties of Arabic Elmadany et al. (2022); Nagoudi et al. (2023b); Alwajih et al. (2024); Abdul-Mageed et al. (2020b, 2021c, 2022, 2023, 2018, c); Keleg et al. (2023); Keleg and Magdy (2023); Zaidan and Callison-Burch (2014); Bouamor et al. (2018). The total number of samples is $122K$. (ii) Synthetically-generated data: As most human-generated dialectal data comes from noisy environments such as social media, it often results in short texts of low quality. Thus, we use Command-R+ to generate paragraph-based synthetic data for Egyptian and Moroccan dialects to improve the performance of our models on dialectal Arabic. We generated $15k$ dialectal instances using the same methodology as our synthetic MSA datasets described above.
Cross-Lingual & Multilingual Datasets. To make our model adapted for cross-lingual and multilingual scenarios, we include the mMARCO dataset, which comprises translations of the MS MARCO dataset into $15$ languages Bonifacio et al. (2021). To ensure that documents correspond accurately to their queries in different languages, we utilize specific IDs. We create $100k$ samples for each cross-lingual pair and shuffle the IDs to prevent repetition, thus guaranteeing that unique data samples are employed for each language. We utilize the MIRACL Zhang et al. (2022), XOR-TyDI Asai et al. (2021), and Mr.TyDi Zhang et al. (2021) datasets as our crosslingual and multilingual resources.
3.2 Training Strategy
For Swan, we consider two models: Swan-Small and Swan-Large. The choice of training two models with different sizes, Swan-Small and Swan-Large, is driven by the need to balance performance and computational efficiency. Swan-Small is designed to cater to scenarios where lower computational resources are available or when a lightweight model is preferred for deployment on edge devices. In contrast, Swan-Large is intended for settings where achieving state-of-the-art performance is paramount, leveraging a larger parameter size to better capture the complexities of the Arabic language and its semantics.
Data Preprocessing. We incorporated human-generated and synthetic datasets into our training pipeline to ensure robust performance across various dialects and cultural contexts. Initially, the model was trained on MSA datasets and then fine-tuned on dialectal datasets. This two-step approach ensures that both MSA and dialectal varieties are well represented, promoting better generalization across the full spectrum of Arabic language varieties. Our dataset was constructed with a query format, including positive and negative samples.
Swan-Small. Built upon ARBERTv2 Abdul-Mageed et al. (2021b), a powerful BERT-based model for the Arabic language. Here, our model is trained using the InfoNCE loss van den Oord et al. (2019a), which maximizes the similarity between related sentences while minimizing the similarity between unrelated sentences. The model was trained for five epochs on the entire dataset with a learning rate of $5e^{-6}$ and a batch size of 128, incorporating fifteen hard negatives. Swan-Small has $164M$ parameters and a dimension size of 768.
Swan-Large. Swan-Large is based on ArMistral-7B, an in-house further pretrained version of Mistral-7B Jiang et al. (2023) (Further details about ArMistral can be found in Appendix A). To train Swan-Large, we use LoRA Hu et al. (2021) for parameter efficient training and InfoNCE loss van den Oord et al. (2019a); Ma et al. (2023) for optimization. The model was trained for three epochs on the entire dataset with a learning rate of $5e^{-6}$ and a batch size of 128, incorporating seven hard negatives. Swan-Large has $7.2B$ parameters and a dimension size of 4096. More details about data preprocessing and training process can be found in Appendix B.
Inclusion of Hard-Negatives To enhance the model’s performance, it is crucial to use negative documents closely aligned with the query’s context Karpukhin et al. (2020). This method allows us to observe the impact of introducing more challenging or "hard" negatives into the training process. We only generate hard negatives for the Arabic subset of our training data from Section 3.1. We found that using $15$ hard negatives for Swan-small yields the best performance, whereas for bigger model Swan-large, the model overfits a more significant number of hard negatives, and $7$ gives us the best performance. (Please see Appendix H for more details about the hard negatives exploration.)
4 ArabicMTEB Benchmark
Task Datasets Languages Dialects Metric
RTR 36 1 4 nDCG@10
CRTR 12 7 0 nDCG@10
CLF 18 1 6 AP
BTM 11 5 8 F1
RRK 5 2 0 MAP
STS 5 1 3 Spearman Corr
CLR 4 1 0 v-measure
PairCLF 3 1 0 AP
Total 94 9 11
Table 3: Overview of our Tasks in ArabicMTEB. Total represents the unique languages.
In this section, we present ArabicMTEB, a comprehensive Arabic-centric text embedding benchmark designed to evaluate text embeddings across a wide range of tasks and scenarios. ArabicMTEB addresses the limitations of existing benchmarks that either exclude Arabic or lack coverage of diverse Arabic language varieties, dialects, and cultural nuances. Our benchmark includes $94$ datasets spanning $8$ distinct tasks, as summarized in Table 3. Further details about the datasets used in  benchmark  can be found in Appendix C. ArabicMTEB was developed to provide comprehensive coverage of Arabic text embedding capabilities, ensuring the inclusion of MSA and other varieties. It offers diverse task types, such as retrieval, classification, and semantic similarity, to evaluate embeddings holistically across different scenarios by incorporating novel domain-specific, dialectal, and country-level culturally aware datasets, ArabicMTEB represents a more applicable and realistic assessment of Arabic text embeddings.
4.1 Task Categories
ArabicMTEB categorizes evaluation datasets into the following key task categories, with each type providing a unique perspective on the capabilities of text embeddings. The corresponding metadata for each task, covering the considered number of datasets, number of languages, number of dialects, and evaluation metric, is presented in Table 3:
Arabic Text Retrieval. This task uses Arabic queries to retrieve Top-$k$ relevant documents from a large Arabic corpus. ArabicMTEB includes $35$ retrieval datasets such as XPDA Shen et al. (2023) and Dolphin’s long-form QA datasets Nagoudi et al. (2023b). Including these datasets helps evaluate complex information retrieval scenarios in Arabic.
Bitext Mining. This task identifies sentence-level translations between different languages and dialects. ArabicMTEB includes $12$ datasets spanning various language pairs like Arabic to French and English. This task is crucial for understanding text embeddings’ cross-lingual and dialectal translation capabilities.
Cross-Lingual Retrieval. This task uses Arabic queries to retrieve documents in other languages, such as English, German, Spanish, and Chinese. ArabicMTEB employs the mMarco Dev set Bonifacio et al. (2021) and includes $11$ language pairs.
Re-Ranking. This task reorders candidate documents for a query based on embedding similarity scores. ArabicMTEB features five re-ranking datasets such as MIRACL Zhang et al. (2023), enabling the evaluation of embeddings’ ability to refine search results.
Semantic Textual Similarity (STS). STS measures the correlation between the embeddings of two sentences, assessing their semantic similarity. ArabicMTEB includes five STS datasets like STS17 and STS22 Cer et al. (2017), along with two synthetic datasets generated using GPT-4 (Details of the creation of these datasets can be found in Appendix I).
Classification. Classification predicts labels for input texts using text embeddings. ArabicMTEB comprises $18$ multi-domain datasets, from ORCA Elmadany et al. (2022). This task evaluates models’ ability to categorize Arabic text accurately, making it a valuable benchmark for downstream tasks such as sentiment analysis.
Pair Classification. This task predicts the relationship between two sentences based on their embeddings. ArabicMTEB includes three datasets, such as XNLI Conneau et al. (2018).
Clustering. Clustering groups sentences into clusters based on embedding similarity, evaluating unsupervised learning performance. ArabicMTEB includes four clustering datasets, such as Arabic News Articles and stance detection datasets from baly-etal-2018-integrating.
4.2 Dialectal ArabicMTEB
Dialectal ArabicMTEB is a specialized fork of the original ArabicMTEB, focusing exclusively on Arabic dialectal datasets. This extension addresses the unique challenges posed by the significant variations in Arabic dialects across different regions, which have been underrepresented in NLP research. While research on dialectal Arabic text embedding has been limited, dialectal ArabicMTEB fills this gap by providing a comprehensive collection of $19$ datasets specifically curated to evaluate embeddings’ performance on diverse Arabic dialects. These datasets span multiple tasks, offering a robust framework for assessing model performance across various dialectal contexts: Bitext Mining: Eight datasets covering dialects such as Algerian, Egyptian, Jordanian, Lebanese, Moroccan, Saudi, and Yemeni Nagoudi et al. (2022); Bouamor et al. (2014). Retrieval: Five datasets focusing on dialects from Algeria, Egypt, Morocco, and the Gulf regions Nagoudi et al. (2023b). Classification: Five datasets for binary, regional, and country-level dialect identification Abdul-Mageed et al. (2021d, 2024); Elmadany et al. (2022); Abdul-Mageed et al. (2021b); Ahmed et al. (2024). STS: A novel synthetic dataset for Egyptian text similarity generated using Command-R+.
4.3 Domain-specific ArabicMTEB
Arabic Text retrieval tasks are currently trending in real-world applications. They are utilized across multiple fields, including healthcare, finance, and legal sectors. Having specialized evaluation datasets is crucial for building text embeddings tailored to these domains. To address this need, we propose Domain-specific ArabicMTEB, a fork of the larger ArabicMTEB benchmark. Domain-specific ArabicMTEB specializes in news, finance, legal, medical, and general knowledge domains, representing a close approximation to real-world scenarios. The creation of this benchmark involved collecting Arabic documents from these specialized sources and from Arabic Wikipedia. We split and chunked these documents into texts of 1,024 tokens in length. Subsequently, we randomly selected chunks and employed GPT4-Turbo OpenAI et al. (2024) to generate five different styles of queries for each chunk. We filtered out duplicate and repeated queries using GPT3.5 OpenAI et al. (2024) to ensure a high-quality evaluation dataset. Our evaluation data creation pipeline is visualized in Figure 2. The resulting benchmark, which we call ArabicMTEB Lite, contains $10k$ queries and $100k$ documents spanning the domains described above.
Figure 3: Generation pipeline for our Domain Specific ArabicMTEB.
4.4 Cultural ArabicMTEB
To show that our models are culturally aware, we have introduced Cultural ArabicMTEB, a collection of datasets from 20 different Arab countries where we focus on specific cultural aspects like their geography, history, etc. To construct the Cultural ArabicMTEB, we use Arabic Wikipedia as our primary data source. For each included Arab country, we extract articles related to that country from its corresponding Wikipedia portal. The portal covers multiple categories (e.g., geography, economy, history) with subcategories (e.g., local movies, food items). This process resulted in $5K$ to $55K$ articles per country. The next step was to generate retrieval questions and passages for each country. For this, we use GPT-4o-mini to develop, for each passage (from an article), a corresponding question whose specific answer is available within the passage itself. Following the same methodology but applied to Egyptian and Moroccan dialectal versions of Wikipedia, we generate dialectal queries and their corresponding passages using Command-R+. Cultural ArabicMTEB  contains $1k$ queries and an average of $15k$ documents from various countries as described above.
5 Evaluation
Model Size Dim. RTR STS PairCLF CLF RRK CLR BTM Avg.
arabertv02-base 160M 768 8.62 39.77 66.30 55.77 60.03 41.74 0.70 38.99
CamelBERT 163M 768 9.21 47.69 67.43 55.66 60.20 39.89 1.85 40.28
ARBERTv2 164M 768 15.12 47.88 68.87 56.85 62.21 39.25 1.99 41.74
ATM-V2 135M 768 37.45 55.90 70.12 46.42 61.45 32.35 12.98 45.24
text2vec 118M 384 27.69 59.37 71.41 47.94 57.76 37.26 38.32 48.54
LaBSE 471M 768 34.98 54.15 70.60 49.57 62.17 41.42 33.28 49.45
Me5-small 118M 384 55.14 56.73 73.97 50.85 67.92 42.37 38.47 55.06
Me5-base 278M 768 56.91 57.99 74.30 52.30 69.07 42.56 33.90 55.29
Swan-Small 164M 768 58.42 59.34 74.93 57.34 68.43 40.43 42.45 57.33
e5-mistral-7b 7110M 4096 56.34 57.02 70.24 53.21 66.24 39.44 70.5 59.00
Me5-large 560M 1024 64.01 59.45 75.06 53.43 70.79 42.49 66.33 61.65
Swan-Large 7230M 4096 65.63 59.10 75.62 54.89 69.42 41.24 71.24 62.45
Table 4: Overall ArabicMTEB results
Model RTR STS CLF BTM Avg.
arabertv02-base 8.67 41.64 47.97 0.99 24.82
MARBERT 5.45 50.06 53.46 2.34 27.83
ARBERTv2 7.52 49.36 54.31 2.51 28.43
CamelBERT 6.92 59.48 50.69 2.65 29.93
AlcLaM 8.56 50.90 54.74 7.54 30.44
ATM-V2 36.23 74.13 34.39 11.67 39.10
Me5-base 61.60 74.84 34.87 3.30 43.65
Me5-small 57.61 76.35 34.78 12.35 45.27
Me5-large 66.88 77.02 35.47 51.08 57.61
e5-mistral-7b 72.35 77.37 35.91 57.62 60.81
Swan-Small 63.16 76.57 54.52 59.38 63.41
Swan-Large 77.03 79.22 53.46 72.10 70.45
Table 5: Dialectal ArabicMTEB results.
Model News Legal Medical Finance Wikipedia Avg Cost $\downarrow$
Swan-Large 90.42 89.96 81.64 57.34 93.10 82.49 0.75\$
Openai-3-large 88.1 89.68 80.24 61.46 91.52 82.20 9.88\$
Cohere-v3.0 85.23 86.52 63.27 42.80 90.96 73.76 7.54\$
Swan-Small 81.55 78.86 70.97 42.48 80.46 70.86 0.44\$
Openai-3-small 71.42 85.23 71.50 32.90 82.20 68.65 3.75\$
Cohere-light-v3.0 70.32 86.83 67.68 22.68 90.34 67.57 2.55\$
Openai-ada-002 65.34 81.83 71.76 39.62 76.79 67.07 1.66\$
Table 6: Domain-Specific ArabicMTEB results.
Model MSA-Culture Egyptian-DIA Morocco-DIA Avg.
Swan-Large 82.19 83.55 65.35 77.03
Cohere-v3.0 81.86 82.90 65.23 76.66
OpenAI-3-large 81.49 78.45 64.90 74.95
Cohere-light-v3.0 80.75 64.82 56.84 67.47
Me5-large 78.65 61.34 60.66 66.88
OpenAI-3-Small 74.55 65.89 54.13 64.86
Swan-Small 75.56 60.35 53.56 63.16
Me5-base 74.56 56.34 53.91 61.60
Me5-small 73.81 53.56 45.45 57.61
ATM-V2 63.78 23.45 21.45 36.23
ARBERTv2 9.34 8.55 4.67 7.52
MARBERT 2.73 0.44 0.19 1.12
Table 7: Cultural ArabicMTEB results.
Model ArRTR DOM-RTR DIA-RTR STS PairCLF CLF RRK CLK BTM Avg.
Swan-Small 15.12 8.46 7.52 37.88 62.87 56.85 62.21 39.25 1.99 32.46
     + Arabic 28.39 39.34 15.23 41.49 70.25 51.89 68.57 39.12 18.74 41.45
     + Synthetic-MSA 31.07 40.45 53.45 55.78 74.23 54.27 68.88 39.43 18.19 48.42
     + Synthetic-DOM 32.01 49.02 49.34 52.90 75.45 54.43 67.45 40.56 17.35 48.72
     + Synthetic-DIA 31.20 38.66 59.43 51.23 72.86 57.56 66.67 37.34 19.90 48.32
Swan-Large 44.46 64.52 66.23 48.63 72.34 50.43 69.39 38.28 44.20 55.39
     + Arabic 54.53 66.43 70.34 52.93 75.24 52.54 70.49 40.21 48.35 59.01
     + Synthetic-MSA 56.34 67.90 72.89 57.89 76.90 50.21 70.92 41.76 62.34 61.91
     + Synthetic-DOM 58.42 76.54 71.65 55.92 75.19 50.19 70.21 39.33 51.23 60.96
     + Synthetic-DIA 57.09 65.06 77.03 56.90 76.42 54.89 69.32 39.41 65.56 62.41
Table 8: The impact of Synthetic Data on Swan performance. ArRTR: Arabic retrieval, DOM-RTR: Domain-specific retrieval, and DIA-RTR: Dialectal Retrieval
We evaluate the performance of our models, Swan-Small and Swan-Large, across the multiple proposed ArabicMTEB benchmarks and compare them with existing state-of-the-art models, including MARBERT Abdul-Mageed et al. (2020a), ARBERTv2 Elmadany et al. (2022), CamelBERT Inoue et al. (2021), multilingual E5 models Wang et al. (2024b), and Arabic-triplet-Matryoshka-V2 (ATM-V2) Nacar and Koubaa (2024). The evaluation results encompasses overall ArabicMTEB  (Table 4), dialectal ArabicMTEB (Table 5), domain-specific ArabicMTEB tasks (Table 6), and cultural ArabicMTEB (Table 7). In these tables, the tasks will be referred to as RTR: Retrieval, STS: Semantic Textual Similarity, PairCLF: Pair Classification, CLF: Classification, CLR: Clustering, RRK: Reranking, and BTM: BiText Mining.
ArabicMTEB Results. Table 4 presents the overall results of our models on the ArabicMTEB benchmark. Our models demonstrate top-tier performance across a variety of NLP tasks. Swan-Small achieves an average score of 57.33, surpassing its main competitors, Me5-base (55.29) and Me5-small (55.06), by a significant margin. This model performs exceptionally well in retrieval (58.42), classification (57.34), and pair classification (74.93), outperforming ATM-V2, which only scores 45.24 on average. Similarly, Swan-Large sets a new state-of-the-art performance with an average score of 62.45, beating Me5-large (61.65) and even the massive e5-mistral-7b model (59.00). The model excels particularly in retrieval (65.63), classification (54.89), and bitext mining (71.24), indicating its robustness across both cross-lingual and Arabic-centric tasks. These results validate our training strategy of using diverse training data covering multiple languages, where Swan-Large outperforms its counterparts by more than five points in cross-lingual tasks such as bitext mining.
Dialectal ArabicMTEB Results. Table 5 shows the dialectal ArabicMTEBresults. Swan-Small scores an average of 63.41, considerably higher than Me5-small (45.27) and AlcLaM (30.44), showing strong performance across retrieval (63.16) and classification (54.52). Swan-Large achieves an impressive average score of 70.45, leading all tasks and outperforming the e5-mistral-7b model, which scores 60.81. The standout result is in bitext mining, which achieves 72.10, showcasing a substantial 14-point improvement over AlcLaM (59.38). Our models’ significant advantage in dialectal retrieval and bitext mining is their unique training with a combination of synthetic and human-generated dialectal datasets, which is absent in many competitive models.
Domain-Specific ArabicMTEB Results. As seen from Table 6, Swan-Small performs exceptionally well, with an average score of 70.86, surpassing OpenAI’s text-embedding-3-small model (68.65) and Cohere-light-v3.0 (67.57). Its best performance is in the legal domain, where it scores 78.86. Swan-Large sets a new standard in domain-specific tasks, scoring 82.49 on average, surpassing OpenAI’s text-embedding-3-large (82.20) and Cohere’s multilingual model (73.76). The model excels particularly in the news domain (90.42), medical (81.64), and Wikipedia (93.10), indicating its superior generalization across varied Arabic domains. Moreover, the cost-effectiveness of our models is evident: using Swan-Large costs only $0.75$ for 10k documents compared to $9.88$ for OpenAI’s model, making it a more efficient solution for large-scale deployments.
Cultural ArabicMTEB Results. Cultural ArabicMTEBis designed to capture culturally sensitive aspects of the Arabic language, such as regional dialects, local idiomatic expressions, and culturally specific knowledge. We generated queries from country-specific Wikipedia articles, including questions about local cuisine, traditional practices, and historical events, which challenge the models to capture more than just linguistic information. For example, Swan-Large achieved the highest performance on tasks related to Egyptian cultural queries, outperforming other models on retrieval tasks by 1.5%. However, we observed slightly lower performance on Moroccan dialect queries, where cultural nuances (such as regional vocabulary) presented a more significant challenge.
Synthetic Data Analysis We systematically analyze the impact of synthetic data on the performance of Swan-Small and Swan-Large using different combinations of training datasets. Table 8 presents the results for the base models, models trained with additional human-generated Arabic data, and models enhanced using synthetic subsets such as MSA, domain-specific, and dialectal data. When comparing the initial Swan-Small (average score of 32.46) to its version trained with synthetic MSA data, we observe a significant increase in average performance to 48.42, representing an improvement of more than 16 points. Similarly, Swan-Large benefits from a 6.52-point boost in average performance (from 55.39 to 61.91) with the inclusion of synthetic MSA data. Inclusion of Synthetic data in Although synthetic data significantly enhances model performance, it can introduce biases due to the reliance on specific patterns in the generated content. We ensured our synthetic data generation diversity by varying the data sources and generating dialectal data for multiple regions, including Egypt, Morocco, and the Gulf states, to mitigate this. We also analysed our models by examining whether MSA data received higher accuracies in retrieval tasks in Table 8. Further, our synthetic data generation pipeline was subjected to human verification for correctness and balance across cultural contexts. Although synthetic data significantly enhances model performance, it can introduce biases due to the reliance on specific patterns in the generated content. We ensured our synthetic data generation diversity by varying the data sources and generating dialectal data for multiple regions, including Egypt, Morocco, and the Gulf states, to mitigate this. We also analysed our models by examining whether MSA data received higher accuracies in retrieval tasks in Table 8. Further, our synthetic data generation pipeline was subjected to human verification for correctness and balance across cultural contexts.
6 Conclusion
In this paper, we introduced Swan-Small and Swan-Large, along with the comprehensive ArabicMTEB  benchmark for evaluating Arabic text embeddings. Our models demonstrate outstanding performance, benefiting from the strategic use of hard negatives and synthetic data in training. The evaluation across multiple benchmarks demonstrates that both Swan-Small and Swan-Large set new standards in Arabic-centric NLP tasks. They outperform existing state-of-the-art models in both cross-lingual and Arabic-specific tasks while being cost-effective and capable of understanding cultural context—making them ideal for real-world applications in diverse Arabic language settings.
7 Limitations
While the development of the Swan  models and the introduction of the ArabicMTEB benchmark mark significant advancements in Arabic text embeddings, there are some limitations to consider.
The societal implications of deploying dialect-aware models, such as Swan, require careful consideration. While these models can bridge gaps in NLP for Arabic-speaking regions, there is a risk of inadvertently reinforcing biases or language hierarchies, particularly in areas where particular dialects are stigmatized or underrepresented. For instance, users in communities with dialects associated with lower socioeconomic status may feel marginalized if their dialect is not adequately supported. To mitigate these concerns, we have prioritized the inclusion of low-resource dialects and ensured that our synthetic data generation pipeline accounts for dialectal diversity. Additionally, future versions of our models will include further dialectal balancing, specifically focusing on underrepresented communities.
8 Ethical Statement
All research and development activities for the Swan models and ArabicMTEB benchmark were conducted with a commitment to ethical standards. Data collection and usage adhered to privacy and confidentiality norms, ensuring no sensitive information was utilized without proper anonymization and consent. We acknowledge the potential biases introduced by synthetic data and have taken steps to mitigate these through diverse data sources and rigorous evaluation.
Appendix A ArMistral Training
ArMistral, is an autoregressive pretrained language model based on Mistral-7B.
Pretraining data We further pretrain it on a large and diverse Arabic dataset, including all categories of Arabic, namely Classical Arabic (CA), Dialectal Arabic (DA), and MSA. This data is aggregated from various sources: AraNewsv2 Nagoudi et al. (2020), El-Khair El-Khair (2016), Gigaword, [^2] LDC Catalog Link OSCAR Suárez et al. (2019), OSIAN Zeroual et al. (2019), 101 Billion arabic words Aloui et al. (2024a), Wikipedia Arabic, and Hindawi Books. [^3] OpenITI corpus (v1.6). We also derived ArabicWeb22 (A) and (B) from the open source Arabic text 2022.444ArabicText-2022 data This pretraining dataset was cleaned, filtered and deduplicated using Bhatia (2023). We have also ensured that the model is pretrained in multiple domains, enhancing its results as seen in Table 9.
Instruction Finetuning. To enhance the capabilities of our ArMistral, we instruct-tuning it on three datasets: Alpaca-GPT4, Evol-instruct, and ShareGPT extracted from MultilingualSIFT datasets Chen et al. (2023).
Alignment Dataset We collected an alignment dataset from Quora and Mawdoo websites and then we took the gold answers as the choosen and we generated the rejected using AceGPT-7B Huang et al. (2024).
Results
Model ARC Hellaswag Exams MMLU Truthfulqa ACVA AlGhafa Average
ArMistral-7B-Chat 43.20 55.53 45.54 43.50 52.44 77.06 35.57 50.41
Jais-13b-chat 41.10 57.70 46.74 42.80 47.48 72.56 34.42 48.97
AceGPT-13B-chat 43.80 52.70 42.09 41.10 49.96 78.42 31.95 48.57
AceGPT-13B-base 39.90 51.30 39.48 40.50 46.73 75.29 30.37 46.22
AraLLama-7B-Chat 39.45 50.23 38.24 41.03 50.44 70.45 32.54 46.05
ArMistral-7B-Base 41.50 52.50 38.92 37.50 51.27 69.64 30.24 45.94
Jais-13b-base 39.60 50.30 39.29 36.90 50.59 68.09 30.07 44.98
AceGPT-7B-chat 38.50 49.80 37.62 34.30 49.85 71.81 31.83 44.81
AraLLama-7B-Base 38.40 50.12 38.43 40.23 45.32 69.42 31.52 44.78
AceGPT-7B-base 37.50 48.90 35.75 29.70 43.04 68.96 33.11 42.42
Table 9: Comparison of ArMistral with other Arabic LLMs
As seen from Table 9, Our ArMistral-Chat model outperforms all existing Arabic LLMs.
Appendix B Training methodology
Given a relevant query-document pair $(q^{+},d^{+})$, we modify the query by appending an instructional template to it. This process transforms the original query $q^{+}$ into a new form $q^{+}_{\text{inst}}$ as defined below:
$q^{+}_{\text{inst}}=\text{Instruction: \{task\_instruction\} Query:}\{q^{+}\}$
Here, “{task_instruction}” refers to a one-sentence description of the embedding task taken from Table 11, which outlines the instructions for different tasks. Using a pretrained large language model (LLM), we append a [EOS] token at the end of both the modified query and the document. These are then input into the LLM to extract embeddings $\mathbf{h}_{q^{+}_{\text{inst}}}$ and $\mathbf{h}_{d^{+}}$ from the vector at the last [EOS] layer. The training of the embedding model is conducted using the InfoNCE loss function van den Oord et al. (2019b), which is widely recognized for its effectiveness in learning high-quality embeddings. The objective is minimized using the following formulation:
$\min\left(-\log\frac{\phi(q^{+}_{\text{inst}},d^{+})}{\phi(q^{+}_{\text{inst}},d^{+})+\sum_{n_{i}\in\mathbb{N}}\phi(q^{+}_{\text{inst}},n_{i})}\right)$
In the equation above, $\mathbb{N}$ denotes the set of negative samples, and $\phi(q,d)$ is the similarity scoring function between a query $q$ and a document $d$.
Appendix C Datasets overview
The table 10 provides a comprehensive summary of the various datasets utilized in the study. It categorizes datasets based on their type, such as Reranking, Bitext Mining, Retrieval, Crosslingual Retrieval, STS, Pair Classification, Clustering, and Classification. Each entry specifies the dataset name, language, citation, and category, reflecting the diversity and scope of data sources for evaluating the model’s performance across different tasks and linguistic contexts.
Appendix D Polydedupe: versatile cleaning Pipeline
PolyDeDupe is a Python package designed for efficient and effective data deduplication across over 100 languages. It supports syntactic and semantic deduplication, making it a versatile tool for high-quality data preprocessing in NLP tasks. Key features include customizable Jaccard similarity thresholds, a performance speed twice that of other tools like SlimPajama, and support for deduplicating instruction tuning data. It can be easily installed via pip to deduplicate datasets, display original and filtered dataset sizes, and identify duplicate clusters. Supported languages span Western, Central, and Eastern European languages, Slavic languages using Cyrillic script, Greek, various Arabic and Devanagari script languages, and more.
Task Dataset Type Language Citation Size
BitextMining Darija S2S Moroccan Arabic Dialect to English Nagoudi et al. (2023b) 2000
BitextMining Narabizi S2S Arabizi to French Nagoudi et al. (2023b) 144
BitextMining Mt_en2ar S2S English to MSA Nagoudi et al. (2023b) 4000
BitextMining Mt_fr2ar S2S French to MSA Nagoudi et al. (2023b) 4000
BitextMining Mt_es2ar S2S Spanish to MSA Nagoudi et al. (2023b) 4000
BitextMining Mt_ru2ar S2S Russian to MSA Nagoudi et al. (2023b) 4000
BitextMining Cs_dz_fr S2S Algerian Arabic Dialect to French Nagoudi et al. (2023b) 200
BitextMining Cs_eg_en S2S Egyptian Arabic Dialect to English Nagoudi et al. (2023b) 200
BitextMining Cs_jo_en S2S Jordanian Arabic to English Nagoudi et al. (2023b) 200
BitextMining Cs_ma_fr S2S Moroccan Arabic to French Nagoudi et al. (2023b) 200
BitextMining Cs_ps_en S2S Palestinian Arabic to English Nagoudi et al. (2023b) 200
BitextMining Cs_ye_en S2S Yemeni Arabic to English Nagoudi et al. (2023b) 200
Classification MassiveIntent S2S Multilingual (Arabic subset) FitzGerald et al. (2022) 100
Classification MassiveScenario S2S Multilingual (Arabic subset) FitzGerald et al. (2022) 100
Classification OrcaSentiment S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaDialect_region S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaDialect_binary S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaDialect_country S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaAns_claim S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaMachine_generation S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaAge S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaGender S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaAdult S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaDangerous S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaEmotion S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaHate_speech S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaOffensive S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaIrony S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaSarcasm S2S Arabic Elmadany et al. (2022) 5000
Classification OrcaAbusive S2S Arabic Elmadany et al. (2022) 5000
Clustering Arabic_news P2P Arabic Our Paper 2500
Clustering Arabic_topic S2S Arabic Our Paper 30
Clustering Arabic_baly_stance P2P Arabic Elmadany et al. (2022) 1000
Clustering Arabic_baly_stance S2S Arabic Elmadany et al. (2022) 100
PairClassification Arabic_xnli S2S Arabic Our Paper 538
PairClassification Arabic_sts S2S Arabic Our Paper 1256
PairClassification Arabic_mq2q S2S Arabic Our Paper 244
Reranking Miracl_ar S2P Multilingual (Arabic subset) Zhang et al. (2023) 750
Reranking Mmarco_arabic S2P Arabic Our Paper 3000
Reranking MedicalQA_arabic S2P Arabic Our Paper 4350
Reranking Mmarco_en2ar S2P English to MSA Our Paper 500
Reranking Mmarco_ar2en S2P MSA to English Our Paper 500
Retrieval MultiLongDoc S2P Multilingual (Arabic subset) MDQA
Retrieval XPQA S2S Multilingual (Arabic subset) XPQA
Retrieval Mintaka S2S Multilingual (Arabic subset) Mintaka
Retrieval Lareqa S2P Arabic Nagoudi et al. (2023b) 220
Retrieval Dawqs S2S Arabic Nagoudi et al. (2023b) 318
Retrieval Exams S2S Arabic Nagoudi et al. (2023b) 2600
Retrieval Mkqa S2S Arabic Nagoudi et al. (2023b) 340
Retrieval Mlqa S2S Arabic Nagoudi et al. (2023b) 517
Retrieval Arcd S2S Arabic Nagoudi et al. (2023b) 693
Retrieval Tydiqa S2S Arabic Nagoudi et al. (2023b) 5700
Retrieval Xsquad S2S Arabic Nagoudi et al. (2023b) 5700
Retrieval Crosslingual_ar2de S2P MSA to German Our Paper 1831
Retrieval Crosslingual_ar2en S2P MSA to English Our Paper 1831
Retrieval Crosslingual_ar2es S2P MSA to Spanish Our Paper 1831
Retrieval Crosslingual_ar2hi S2P MSA to Hindi Our Paper 1831
Retrieval Crosslingual_ar2vi S2P MSA to Vietnamese Our Paper 1831
Retrieval Crosslingual_ar2zh S2P MSA to Chinese Our Paper 1831
Retrieval Crosslingual_de2ar S2P German to MSA Our Paper 1831
Retrieval Crosslingual_en2ar S2P English to MSA Our Paper 1831
Retrieval Crosslingual_es2ar S2P Spanish to MSA Our Paper 1831
Retrieval Crosslingual_hi2ar S2P Hindi to MSA Our Paper 1831
Retrieval Crosslingual_vi2ar S2P Vietnamese to MSA Our Paper 1831
Retrieval Crosslingual_zh2ar S2P Chinese to MSA Our Paper 1912
Retrieval MoroccoCultural S2P Arabic Our Paper 100
Retrieval SyriaCultural S2P Arabic Our Paper 100
Retrieval LibyaCultural S2P Arabic Our Paper 100
Retrieval LebanonCultural S2P Arabic Our Paper 100
Retrieval QatarCultural S2P Arabic Our Paper 100
Retrieval SudanCultural S2P Arabic Our Paper 100
Retrieval AlgeriaCultural S2P Arabic Our Paper 100
Retrieval MauritaniaCultural S2P Arabic Our Paper 100
Retrieval TunisiaCultural S2P Arabic Our Paper 100
Retrieval IraqCultural S2P Arabic Our Paper 100
Retrieval EgyptCultural S2P Arabic Our Paper 100
Retrieval SomaliaCultural S2P Arabic Our Paper 100
Retrieval UAE_Cultural S2P Arabic Our Paper 100
Retrieval OmanCultural S2P Arabic Our Paper 100
Retrieval KuwaitCultural S2P Arabic Our Paper 100
Retrieval BahrainCultural S2P Arabic Our Paper 100
Retrieval Saudi_ArabiaCultural S2P Arabic Our Paper 100
Retrieval JordanCultural S2P Arabic Our Paper 100
Retrieval PalestineCultural S2P Arabic Our Paper 100
Retrieval YemenCultural S2P Arabic Our Paper 100
Retrieval MoroccoDIA S2P Moroccan Arabic Dialect Our Paper 100
Retrieval EgyptDIA S2P Egyptian Arabic Dialect Our Paper 100
Retrieval NewsDomainSpecific S2P Arabic Our Paper 1000
Retrieval LegalDomainSpecific S2P Arabic Our Paper 1000
Retrieval MedicalDomainSpecific S2P Arabic Our Paper 1000
Retrieval FinanceDomainSpecific S2P Arabic Our Paper 1000
Retrieval WikipediaDomainSpecific S2P Arabic Our Paper 1000
STS STS17 S2S Arabic Cer et al. (2017) 8060
STS STS22 P2P Arabic Semenov et al. (2023) 500
STS Arabic_sts S2S Arabic Our Paper 750
STS Arabic_stsb_multi_dialect S2S Arabic Dialectal Our Paper 1500
STS Arabic_sts P2P Arabic Our Paper 500
Table 10: Benchmark Datasets Overview. Abbreviations: S2S = Sentence to Sentence, S2P = Sentence to Paragraph, P2P = Paragraph to Paragraph.
Appendix E Prompts for evaluation
Table 11 provides an overview of the prompts used for evaluating various tasks. It includes instructions for Reranking, Bitext Mining, Retrieval, Crosslingual Retrieval, Semantic Textual Similarity (STS), Pair Classification, Clustering, and Classification. Each entry outlines the specific task and the corresponding instruction used to guide the model’s evaluation process.
Task Instructions
Reranking Given an Arabic search query, retrieve web passages that answer the question in {Lang}. Query:{query}.
BitextMining Retrieve parallel sentences in {Lang}.
Retrieval Given an Arabic search query, retrieve web passages that answer the question. Query:{query}.
Crosslingual Retrieval Given an Arabic search query, retrieve web passages that answer the question in {Lang}. Query:{query}.
STS Retrieve semantically similar text. Text: {text}.
Pair Classification Retrieve texts that are semantically similar to the given text. Text: {text}.
Clustering Identify the topic or theme of the given news article. Article:{article}.
Classification Classify the text into the given categories {options}.
Table 11: Prompts used for evaluation.
Appendix F Full Leaderboard
Table 12 presents the performance comparison of various models on different tasks within the ArMTEB benchmark. It includes metrics for Retrieval, Semantic Textual Similarity (STS), Pair Classification (PairCLF), Classification (CLF), Re-ranking, Clustering, and Bitext Mining (BTM). The table lists each model, its dimensionality, and the scores for each task, along with an overall average score. The results highlight the strengths and weaknesses of each model across a range of tasks, providing a comprehensive overview of their performance.
Model Dim. Retrieval STS PairCLF CLF Re-rank Cluster BTM Avg
Number of datasets 23 5 3 18 5 4 12 70
Swan-Large 4096 65.63 59.10 75.62 52.55 69.42 41.24 71.24 62.11
multilingual-e5-large 1024 64.01 59.45 75.06 53.43 70.79 42.49 66.33 61.65
e5-mistral-7b-instruct 4096 56.34 57.02 70.24 53.21 66.24 39.44 70.50 59.00
Swan-Base 768 58.42 58.44 74.93 57.34 68.43 40.43 42.45 57.21
multilingual-e5-base 768 56.91 57.99 74.30 52.30 69.07 42.56 33.90 55.29
multilingual-e5-small 384 55.14 56.73 73.97 50.85 67.92 42.37 38.47 55.06
LaBSE 768 34.98 54.15 70.60 49.57 62.17 41.42 33.28 49.45
text2vec-base 384 27.69 59.37 71.41 47.94 57.76 37.26 38.32 48.54
ARBERTv2 768 15.12 37.88 62.87 56.85 62.21 39.25 1.99 39.45
CamelBERT-msa 768 9.21 47.69 67.43 55.77 60.20 39.89 1.85 40.29
arabertv02-large 1024 7.34 34.26 63.63 54.32 56.71 37.26 10.97 37.78
arabertv02-base 768 8.62 39.77 66.30 55.77 60.03 41.74 0.70 38.99
CamelBERT-mix 768 7.19 46.47 67.23 56.68 57.50 38.72 0.41 39.17
MARBERTv2 768 5.88 45.21 70.89 54.89 58.64 40.81 0.45 39.54
ARBERT 768 8.07 29.89 61.86 56.92 61.09 37.10 2.28 36.74
CamelBERT-da 768 4.07 41.05 65.82 53.75 54.44 37.63 0.31 36.72
MARBERT 768 2.22 40.62 66.46 54.35 53.09 36.33 0.40 36.21
CamelBERT-ca 768 2.74 36.49 62.26 46.26 51.34 35.77 0.09 33.56
Table 12: ArMTEB Results.
Appendix G Inference Latency.
Figure 4: Latency vs Performance.
Inference latency is very critical in deploying machine learning models, especially in real-time applications with crucial response time. It refers to the time taken by a model to predict received input. In the context of text embedding models such as Swan-Base and Swan-Large, lower latency is particularly valuable for user-facing services that rely on fast processing of natural language input, such as chatbots and search engines. From Figure 4, we find that Swan-Large, despite its larger size indicated by a larger bubble, has optimized inference times due to architectural efficiencies, and Swan-Base strikes the perfect balance between size, performance, and latency. We compare the performance of the models from Table 4.
Appendix H Impact of Hard Negatives
Model (HN) 1 3 7 15 31
Swan-Small 48.84 52.19 54.13 56.25 51.93
Swan-Large 59.48 59.35 60.42 59.44 59.83
Table 13: Impact of number of Hard Negatives (HN).
Hard negatives are challenging examples that are nearly correct but ultimately incorrect, forcing the model to learn more nuanced distinctions between the different classes. The process involves converting all documents into a vector form within the embedding space. Subsequently, these document embeddings are compared using the cosine similarity score to establish their relevance to the query. Once all documents are scored, they are sorted by their similarity to the query. The top-ranked document is typically the positive example, while the rest are potential negatives. Our experiments assess the impact of varying the hard negatives used while training our models, Swan-Large and Swan-Base. We train each model with different quantities of hard negatives. Namely, we experiment with using 1, 3, 7, 15, and 31 hard negatives per training instance. Swan-Base reaches its highest performance at 56.25 with 15 hard negatives. This model shows a general upward trend in performance as the number of hard negatives increases, peaking at 15 but then declining slightly when the number increased to 31. This pattern suggests that while additional hard negatives initially provide beneficial learning challenges, there can be a point of diminishing returns where too much complexity hinders further learning. Swan-Large shows a peak in performance with $60.42$ when trained with seven hard negatives, indicating an optimal level of challenge that enhances learning without overwhelming the model. Interestingly, further increases in hard negatives do not improve performance, suggesting a threshold beyond which additional complexity does not translate to better learning outcomes.
Appendix I STS Dataset Creation:
The Arabic Semantic Textual Similarity (Arabic-STS) datasets was developed to facilitate research in semantic similarity for the Arabic language. The dataset is derived from the Arabic Billion Words Aloui et al. (2024b) corpus, which serves as a foundation for extracting a diverse collection of sentence pairs. Each pair is annotated with a similarity score that captures the degree of semantic equivalence between the sentences. The dataset generation process was guided by the capabilities of the GPT-4 model developed by OpenAI, ensuring that the resulting sentence pairs are of high quality and reflect nuanced linguistic characteristics. The creation involved several steps, including selecting representative sentences from the corpus, generating semantically varied sentence pairs, and annotating similarity scores using both automated methods and human reviewers to maintain consistency and reliability.
Appendix J Country level Cultural Evaluation
Model Swan-Large Me5-large Cohere-light-v3.0 Swan-Base OpenAI-3-large Cohere-v3.0 Me5-small Me5-base ATM-V2 ARBERTv2 MARBERT
Algeria 89.34 93.34 89.44 90.45 86.95 88.99 91.23 90.66 84.99 18.27 1.50
Bahrain 93.71 93.77 93.52 86.48 91.98 92.40 93.08 89.04 90.49 27.48 5.74
Egypt 98.34 94.58 91.37 95.66 91.45 87.81 93.02 91.65 88.45 11.54 1.63
Iraq 92.45 90.90 86.98 88.34 92.43 87.83 89.02 90.78 81.22 17.34 1.92
Jordan 92.34 92.79 90.07 89.70 94.56 91.18 93.67 92.25 87.95 27.46 4.50
Kuwait 93.45 96.34 96.10 90.44 88.53 92.51 96.17 94.94 89.97 36.67 4.92
Lebanon 95.66 93.05 92.38 90.45 90.23 91.04 91.92 92.85 87.14 22.55 1.82
Libya 89.56 88.43 87.27 85.45 89.66 85.75 87.21 85.32 79.95 28.88 2.46
Mauritania 92.44 92.92 92.61 89.45 90.31 92.05 20.99 3.32 0.63 0.50 0.00
Morocco 90.34 85.49 83.19 86.34 83.56 85.47 81.73 86.59 4.75 0.32 0.00
Oman 94.45 94.26 92.37 91.98 92.45 92.61 93.00 93.04 84.21 11.24 3.43
Palestine 90.45 90.67 87.50 91.18 87.45 83.33 85.22 86.49 77.83 27.25 3.63
Qatar 98.79 93.44 91.80 92.35 95.66 89.98 91.20 90.49 85.50 29.15 7.00
Saudi_Arabia 95.34 93.49 92.98 91.47 90.45 92.12 92.72 91.47 86.48 25.06 2.50
Somalia 90.23 94.78 93.67 88.34 89.55 92.30 21.25 2.50 20.81 2.62 0.00
Sudan 92.36 91.99 86.90 90.89 91.45 90.72 89.49 87.60 82.47 24.51 2.50
Syria 91.46 91.83 90.56 90.45 90.56 86.97 88.69 88.75 87.45 13.81 3.63
Tunisia 94.57 94.64 93.46 95.54 85.34 90.92 93.79 92.04 84.40 25.04 4.15
UAE 96.09 95.14 93.41 94.12 97.66 93.53 94.45 91.56 91.79 31.92 2.00
Yemen 92.34 91.24 89.40 92.12 89.54 89.70 88.25 89.89 83.08 5.29 1.29
Avg. 93.19 92.65 90.75 90.56 90.49 89.86 83.81 81.56 73.98 19.34 2.73
Table 14: Country level Cultural evaluation